OpenStudy (anonymous):

what is x(x+6) in expanded form

OpenStudy (zzr0ck3r):

distribute

OpenStudy (anonymous):

x+6x because youre distributing the x. x times x = x and x times 6 = 6x. your new equation would be x+6x but the solution would be 7x

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

\(\large x(x+6) = x^2+6x\)

OpenStudy (zzr0ck3r):

\(x*x=x^2\) @ninamarie

OpenStudy (anonymous):

youre right sorry lol havnt done this in a while